# RAZOR: Sharpening Knowledge by Cutting Bias with Unsupervised Text Rewriting

### v1

@@ 4.1 Discussion
We train BERT and RoBERTa classifiers on the original FEVER dataset to represent our baseline (hereafter Original). Then, for each SoTA method, we train the two classifiers on the modified version of FEVER according to the proposed algorithms. Notice that RAZOR and other SoTA methods only rewrite the training set of the dataset. The test set remains untouched. We follow this strategy to assess the impact of the rewriting strategy w.r.t. the test set’s original peculiarities, which might still contain shortcuts. We expect that if the rewriting strategy truly alleviates the shortcut problem in the training set, then in the test set, we will see an increase in performance since the false negative/positive ratio should decrease [^6] As per Def. 1, a shortcut causes classifiers to have false positives/negatives.. Table 1 shows this effect on the two classifiers mentioned above for RAZOR and SoTA methods. Notice how, although the second-best after CrossAug, RAZOR effectively removes the shortcut effect on both classifiers w.r.t. the Original baseline. We acknowledge that RAZOR is behind CrossAug in terms of performance because the latter enriches the dataset with contrastive – i.e., positive-negative – claim-evidence pairs, where the evidence is rewritten differently from RAZOR, which only operates at the claim level.
Nevertheless, we show how RAZOR is more robust than CrossAug on the FEVER-Adversarial dataset. [^7] The training set of FEVER and FEVER-Adversarial is the same. The test sets are different. The sentences in the test set of FEVER-Adversarial do not contain shortcuts. Hence, we use the adversarial test set to demonstrate RAZOR’s debiasing abilities. Notice how CrossAug’s performance is systematically lower than that of RAZOR. Because CrossAug negates the claim and evidence to generate other factual instances, it might happen that this newly generated claim-evidence pair can itself contain shortcuts, exacerbating the classifier’s reasoning capabilities. However, since FEVER’s test set contains shortcuts, CrossAug’s original performances are not faithful to assess the goodness of this debiasing strategy. In particular, this is evident in FEVER-Adversarial, where simple negations of claim-evidence pairs are not optimal. Contrarily, RAZOR successfully reduces the effect of the shortcuts and achieves consistently better results than all SoTa methods.
@@
To validate the generalizability of our method, we rely on MNLI and SNLI for a multi-class classification test. Similarly to the above process, we fine-tuned BERT and RoBERTa as a baseline to determine whether the logical relationship between two sentences is entailment, contradiction, or neutral. Subsequently, we used RAZOR to rewrite the training sets and achieve improvements in Accuracy and F1 scores for the two classifiers. With the lack of corresponding attacking datasets like FEVER-Adversarial, we exchanged MNLI and SNLI test sets to simulate a real-world data distribution.
Our experiments demonstrate that even widely used datasets exhibit shortcut problems. We observed that by applying RAZOR, the baseline classifiers achieved improvements of approximately 0.1 to 8 points across all datasets and metrics. The most significant improvement is observed in the SNLI-MNLI training-test set combination. This suggests that SNLI may contain more shortcuts than MNLI, resulting in a performance difference of over 20 points for BERT across different test sets. However, after applying RAZOR, a substantial portion of the shortcuts in the data was removed, leading to an increase of 8.14 points in F1 score and 5.63 points in accuracy for BERT. Moreover, we find that RAZOR performs better on BERT than on RoBERTa overall. This is because models with better overall performances are less susceptible to shortcut issues. Therefore, RAZOR demonstrates significant potential for debiasing smaller models.
@@ 5 Conclusion
RAZOR is an effective dataset debiasing method. It outperforms SoTA methods in addressing shortcut problems in fact-checking and natural language inference tasks. RAZOR maintains dataset integrity by generating new sentences that adhere to the original labels and strategically removing the original documents while significantly reducing shortcut-induced biases. This is achieved through an iterative rewriting process, optimized until the KL divergence between the embeddings in the surface feature space is minimized. Our experiments consistently show that RAZOR enhances the performance of classifiers such as BERT and RoBERTa, particularly in challenging adversarial scenarios where traditional methods like CrossAug fail to deliver competitive results. Notably, RAZOR’s superior performance on models like BERT, even compared to more robust models like RoBERTa, underscores its potential as a powerful debiasing tool, particularly for smaller and more resource-constrained models. The implications of our work extend beyond the datasets and models tested, offering a promising approach to improving model reliability across a wide range of machine learning applications. While RAZOR has shown great promise, future work could explore its application to other types of biases or in conjunction with additional debiasing strategies. We believe that RAZOR represents a significant step forward in the ongoing effort to create more fair and accurate machine learning systems, ultimately contributing to more trustworthy AI applications in the real world.
@@ Appendix A Proof of Lemma 1
During training, the model adjusts its parameters to minimize a loss function $\mathcal{L}$, which measures the difference between the model’s predictions and the true labels. The attention scores $\mathbf{z}_{j}=h(g(t_{j}))$ for each token $t_{j}\in d_{i}$ in the sequence are influenced by this optimization process. Because of Eq. (1), the information in $\hat{d}_{i}$ is sufficient for $f_{\theta}$ to make its decision. Now, the attention mechanism in transformers is designed to assign higher weights to more relevant tokens. Given that $\hat{d}_{i}$ contains sufficient information for the prediction, we can safely assume that
@@
Let us consider the left side of the inequality of Eq. 4 – see Eq. 5 – by the triangle inequality we have
@@ Appendix B Additional Experiments
Figure 5: The performance of models trained on 500 samples which randomly selected from FEVER when evaluated on the FEVER-Adverasial test set.
@@
RAZOR does not require large datasets to correctly remove shortcuts in fact-checking datasets.
Fig. 5 presents additional results for our ablation study. We observed that under a small data size, RAZOR still yields noticeable improvements of 2-4 points in Accuracy and F1 score. Furthermore, compared to the embedding-based RAZOR, we analyze that the improvement is attributed solely to eliminating surface shortcuts rather than using GPT-3.5-Turbo.
@@
Table 4 presents the corpus-level BLEU scores (Papineni et al. 2002) for the original training set and the rewritten ones with RAZOR, evaluated on the tests sets of FEVER and FEVER-Adversarial. After rewriting, the BLEU scores between the original data and the Adversarial test set which is more generalizable increased twofold, indicating that RAZOR effectively mitigates biases at the surface feature level.
@@ Appendix C User-Evaluation Survey and Examples for Explanations on Fact-Checking Datasets
Fact verification means utilizing Wikipedia as evidence to evaluate user-generated claims online, determining whether factual evidence supports or refuses these claims. We conducted a user-study evaluation test [^8] We had 10 participants in the survey. on ten randomly selected evidence samples in the FEVER dataset. Each question contained the original claim-evidence pair and the corresponding rewritten claim with RAZOR for the exact same evidence. The tokens on the evidence for the original and rewritten claim are colored according to SHAP – i.e., the bluer the token, the more support for the claim; the redder, the more the refusal. The participants were asked to rate the rationality and comprehensiveness of the explanations on the evidence for the original and rewritten claims on a scale of 1 to 7. The participants were also asked to choose their preference between the explanation of the original claim and the rewritten claim. Here, we want to demonstrate that the debiased model – in this case, RoBERTa w/ RAZOR – is less influenced by shortcuts compared to the original model, and, thus, RAZOR aids in generating more reasonable classification “justifications”. Table 5 illustrates the results of this study. Note that the rewritten claims contain better explanations with a margin of 1.64 and 1.46 in, respectively, reasonableness and comprehensiveness. Moreover, we observe a 79% preference for the explanations on the evidence for the rewritten claim with both refusal and support tokens – i.e., RoBERTa w/ RAZOR - from the participants, leading us to believe that human-centered evaluations add-on to the usefulness of RAZOR besides it having SoTA performances (see Tables 1 and 2).
@@ Appendix D RAZOR Algorithm
Algorithm 1 illustrates the entire process of how RAZOR operates. In the rewriting process – in particular, line 13 – we assume that it implictly checks for consistency (see Eq. (13)). Therefore, $\varphi(d_{i})$ here only contains rewritten documents that agree with the label $\Phi(d_{i})$.
@@
$\sum_{d_{i}\in\mathcal{D}_{0}}\sum_{d_{j}\in\mathcal{D}_{1}}\text{KL}(\hat{g}(d_{i}),\hat{g}(d_{j}))$
is minimized where $\mathcal{D}_{y}=\{d\;|\;d\in\mathcal{D}\;\land\;\Phi(d)=y\}$

### v2

@@ 4.1 Discussion
We train BERT, RoBERTa, and DistilBERT classifiers on the original FEVER dataset to represent our baseline (hereafter Original). Then, for each SoTA method, we train the two classifiers on the modified version of FEVER according to the proposed algorithms. Notice that RAZOR and other SoTA methods only rewrite the training set of the dataset. The test set remains unchanged. We follow this strategy to assess the impact of the rewriting strategy w.r.t. the test set’s original peculiarities, which might still contain shortcuts. We expect that if the rewriting strategy truly alleviates the shortcut problem in the training set, then in the test set, we will see an increase in performance since the false negative/positive ratio should decrease [^6] As per Definition 1, a shortcut causes classifiers to have false positives/negatives.. Table 1 shows this effect on the two classifiers mentioned above for RAZOR and SoTA methods. Notice how, although second-best after CrossAug, RAZOR effectively removes the shortcut effect on both classifiers w.r.t. the original baseline. We acknowledge that RAZOR is behind CrossAug in terms of performance because the latter enriches the dataset with contrastive – i.e., positive-negative – claim-evidence pairs, where the evidence is rewritten differently from RAZOR, which only operates at the claim level. However, we show how RAZOR is more robust than CrossAug in the FEVER-Adversarial dataset. [^7] The training set of FEVER and FEVER-Adversarial is the same. The test sets are different. The sentences in the test set of FEVER-Adversarial do not contain shortcuts. Hence, we use the adversarial test set to demonstrate RAZOR’s debiasing abilities. Notice how CrossAug’s performance is systematically lower than that of RAZOR. Because CrossAug negates the claim and evidence to generate other factual instances, it might happen that this newly generated claim-evidence pair can itself contain shortcuts, exacerbating the classifier’s reasoning capabilities. However, since FEVER’s test set contains shortcuts, CrossAug’s original performances are not faithful to assess the goodness of this debiasing strategy. In particular, this is evident in FEVER-Adversarial, where simple negations of claim-evidence pairs are not optimal. Contrarily, RAZOR reduces shortcuts’ effect and achieves consistently better results than all SoTA methods.
@@
To validate the generalizability of our method, we rely on MNLI and SNLI for a multi-class classification test. Similarly to the above process, we fine-tuned the classifiers as baselines to determine whether the logical relationship between two sentences is entailment, contradiction, or neutral. Subsequently, we used RAZOR to rewrite the training sets and achieve improvements in Accuracy and F1 scores for the two classifiers. With the lack of corresponding attacking datasets like FEVER-Adversarial, we exchanged MNLI and SNLI test sets to simulate a real-world data distribution.
Our experiments demonstrate that even widely used datasets exhibit shortcut problems. We observed that by applying RAZOR, the baseline classifiers achieved improvements of approximately 0.1 to 8 points across all datasets and metrics. The most significant improvement is observed in the SNLI-MNLI training-test set combination. This suggests that SNLI may contain more shortcuts than MNLI, resulting in a performance difference of over 20 points for BERT across different test sets. However, after applying RAZOR, a substantial portion of the shortcuts in the data was removed, leading to an increase of 8.14 points in F1 score and 5.63 points in accuracy for BERT. Moreover, we find that RAZOR performs better on all classifiers overall. This is because models with better overall performances are less susceptible to shortcut issues. Therefore, RAZOR demonstrates significant potential for debiasing smaller models.
@@ 5 Conclusion
RAZOR is an effective dataset debiasing method. It outperforms SoTA methods in addressing shortcut problems in fact-checking and natural language inference tasks. RAZOR maintains dataset integrity by generating new sentences that adhere to the original labels and strategically removing the original documents while significantly reducing shortcut-induced biases. This is achieved through an iterative rewriting process, optimized until the KL divergence between the embeddings in the surface feature space is minimized. Our experiments consistently show that RAZOR enhances the performance of classifiers such as BERT and RoBERTa, particularly in challenging adversarial scenarios where traditional methods like CrossAug fail to deliver competitive results. Notably, RAZOR’s superior performance on models like BERT, even compared to more robust models like RoBERTa, underscores its potential as a powerful debiasing tool, particularly for smaller and more resource-constrained models. The implications of our work extend beyond the datasets and models tested, offering a promising approach to improving model reliability across a wide range of machine-learning applications. While RAZOR has shown great promise, future work could explore its application to other types of biases or in conjunction with additional debiasing strategies. We believe that RAZOR represents a significant step forward in the ongoing effort to create more fair and accurate machine learning systems, ultimately contributing to more trustworthy AI applications in the real world.
@@ Appendix A Proof of Lemma 1
During training, the model adjusts its parameters to minimize a loss function $\mathcal{L}$, which measures the difference between the model’s predictions and the true labels. The attention scores $\mathbf{z}_{j}=h(g(t_{j}))$ for each token $t_{j}\in d_{i}$ in the sequence are influenced by this optimization process. Because of Eq. (1), the $\hat{d}_{i}$ information is sufficient for $f_{\theta}$ to decide. Now, the attention mechanism in transformers is designed to assign higher weights to more relevant tokens. Given that $\hat{d}_{i}$ contains sufficient information for the prediction, we can safely assume that
@@
Let us consider the left side of the inequality of Eq. 4 – see Eq. 5 – by the triangle inequality, we have
@@ Appendix B Additional Experiments
Figure 5: The performance of models trained on 500 samples randomly selected from FEVER when evaluated on the FEVER-Adverasial test set.
@@
RAZOR does not require large datasets to remove shortcuts in fact-checking datasets correctly.
Fig. 5 presents additional results for our ablation study. We observed that under a small data size, RAZOR still yields noticeable improvements of 2-4 points in Accuracy and F1 score. Furthermore, compared to the embedding-based RAZOR, we analyze that the improvement is attributed solely to eliminating surface shortcuts rather than GPT-3.5-Turbo.
@@
Table 3 presents the corpus-level BLEU scores (papineni-etal-2002-bleu) for the original training set and the rewritten ones with RAZOR, evaluated on the tests sets of FEVER and FEVER-Adversarial. After rewriting, the BLEU scores between the original data and the Adversarial test set, which is more generalizable, increased twofold, indicating that RAZOR effectively mitigates biases at the surface feature level.
@@ Appendix C User-Evaluation Survey and Examples for Explanations on Fact-Checking Datasets
Fact verification means utilizing Wikipedia as evidence to evaluate user-generated claims online, determining whether factual evidence supports or refuses these claims. We conducted a user-study evaluation test [^8] We had 10 participants in the survey. on ten randomly selected evidence samples in the FEVER dataset. Each question contained the original claim-evidence pair and the corresponding rewritten claim with RAZOR for the same evidence. The tokens on the evidence for the original and rewritten claim are colored according to SHAP – i.e., the bluer the token, the more support for the claim; the redder, the more the refusal. The participants were asked to rate the rationality and comprehensiveness of the explanations on the evidence for the original and rewritten claims on a scale of 1 to 7. The participants were also asked to choose their preference between the explanation of the original claim and the rewritten claim. Here, we want to demonstrate that the debiased model – in this case, RoBERTa w/ RAZOR – is less influenced by shortcuts compared to the original model, and, thus, RAZOR aids in generating more reasonable classification “justifications”. Table 4 illustrates the results of this study. Note that the rewritten claims contain better explanations with a margin of 1.64 and 1.46 in, respectively, reasonableness and comprehensiveness. Moreover, we observe a 79% preference for the explanations on the evidence for the rewritten claim with both refusal and support tokens – i.e., RoBERTa w/ RAZOR - from the participants, leading us to believe that human-centered evaluations add-on to the usefulness of RAZOR besides it having SoTA performances (see Tables 1 and 2).
@@ Appendix D RAZOR Algorithm
Algorithm 1 illustrates the entire process of how RAZOR operates. In the rewriting process – in particular, line 13 – we assume that it implicitly checks for consistency (see Eq. (13)). Therefore, $\varphi(d_{i})$ here only contains rewritten documents that agree with the label $\Phi(d_{i})$.
@@
$\sum_{d_{i}\in\mathcal{D}_{0}}\sum_{d_{j}\in\mathcal{D}_{1}}cos(\hat{g}(d_{i}),\hat{g}(d_{j}))$
is maximized where $\mathcal{D}_{y}=\{d\;|\;d\in\mathcal{D}\;\land\;\Phi(d)=y\}$